The mass of Earth is 6.0 * 10^24kg. The mass of the moon is 7.4 * 10^22kg. The distance between the Earth and the moon is 3.8 * 10^8m.


(a) Treating the bodies as point masses, calculate the gravitational field strength at a distance of 3.42 × 10^8 m from the Earth, on the line that joins them.


b) Comment on this result; what is the net field strength at that point? (1 mark)


1
Expert's answer
2021-08-17T10:30:12-0400

The field from the earth:


"g_e=\frac{GM_e}{r^2}=0.00342\text{ N/m},\\\space\\
g_m=\frac{GM_m}{(R-r)^2}=0.00342\text{ N/m}."

The gravitational field strength is 0 if we consider 3 significant figures and it's 0.00004 of we consider more.


Add we see, the gravitational influence of the earth is infinitesimal.